# Sample Complexity of Linear Quadratic Regulator Without Initial Stability

### v2

@@ 4. The Inner Loop and Policy Gradient
$Q_{h}(x_{h},-Kx_{h}+\sigma\eta_{h})$
$=$ $x_{h}^{\top}Qx_{h}+(-Kx_{h}+\sigma\eta_{h})^{\top}R(-Kx_{h}+\sigma\eta_{h})+(Ax_{h}+B(-Kx_{h}+\sigma\eta_{h}))^{\top}\widetilde{P}_{h+1}(Ax_{h}+B(-Kx_{h}+\sigma\eta_{h}))$
$\leq$ $\|Q\|C_{m}+\|R\|\|-Kx_{h}+\sigma\eta_{h}\|^{2}+\|\widetilde{P}_{h+1}\|\|Ax_{h}+B(-Kx_{h}+\sigma\eta_{h})\|^{2}.$
@@
(54) $Q_{h}(x_{h},-Kx_{h}+\sigma\eta_{h})$
(55) $\leq$ $\|Q\|C_{m}+2\|R\|(\widetilde{C}_{h}C_{m}+\sigma^{2}\|\eta_{h}\|^{2})+2(\|P_{\max}\|+a)\|A\|^{2}C_{m}+4(\|P_{\max}\|+a)\|B\|^{2}(\widetilde{C}_{h}C_{m}+\sigma^{2}\|\eta_{h}\|^{2}))$
(56) $=$ $C_{m}(\|Q\|+2\|R\|\widetilde{C}_{h})+2C_{m}(\|P_{\max}\|+a)(\|A\|^{2}+2\|B\|^{2}\widetilde{C}_{h}))+2\left(\|R\|+2(\|P_{\max}\|+a)\|B\|^{2}\right)\sigma^{2}\|\eta_{h}\|^{2},$
@@
(57) ^$\|\widehat{\nabla}\widetilde{J}_{h}(K)\|_{F}$
(58) $\leq$ $\frac{1}{\sigma}\Bigg{(}\|Q\|+2\|R\|\widetilde{C}_{h}+2(\|P_{\max}\|+a)(\|A\|^{2}+2\|B\|^{2}\widetilde{C}_{h})\Bigg{)}C_{m}^{3/2}\|\eta_{h}\|$
(59) $+2\left(\|R\|+2(\|P_{\max}\|+a)\|B\|^{2}\right)\sigma C_{m}^{1/2}\|\eta_{h}\|^{3}$
(60) $=$ $\frac{1}{\sigma}\xi_{h,1}\|\eta_{h}\|+\sigma\xi_{h,2}\|\eta_{h}\|^{3}.$
Furthermore, since $\eta_{h}\sim\mathcal{N}(0,I_{m})$ for any $h$, $\|\eta_{h}\|^{2}$ is distributed according to the chi-squared distribution with $m$ degrees of freedom ($\|\eta_{h}\|^{2}\sim\chi^{2}(m)$ for any $h$). Therefore, the standard Laurent-Massart bounds [10] suggest that for arbitrary $y>0$, we have that
(61) ${\mathbb{P}}\{\|\eta_{h}\|^{2}\geq m+2\sqrt{my}+2y\}\leq e^{-y}.$
Now if we take $y=m\log\frac{1}{\zeta}$, since $\zeta\in(0,1/e)$ by our assumption, it holds that $y=m\log\frac{1}{\zeta}\geq m$. Thus
${\mathbb{P}}\{\|\eta_{h}\|^{2}\geq 5y\}$$\leq{\mathbb{P}}\{\|\eta_{h}\|^{2}\geq m+2\sqrt{my}+2y\}$
$\leq e^{-y},$
which after substituting $y$ with its value $m\log\frac{1}{\zeta}$ gives
${\mathbb{P}}\{\|\eta_{h}\|^{2}\geq 5m\log\frac{1}{\zeta}\}\leq e^{-m\log\frac{1}{\zeta}}=\zeta^{m}\leq\zeta.$
As a result, we have $\|\eta_{h}\|\leq 5^{1/2}m^{1/2}(\log\frac{1}{\zeta})^{1/2}$ and consequently
$\|\eta_{h}\|^{3}\leq 5^{3/2}m^{3/2}(\log\frac{1}{\zeta})^{3/2}$
with probability at least $1-\zeta$, which after applying on (60) yields
^$\|\widehat{\nabla}\widetilde{J}_{h}(K)\|_{F}$ $\leq\frac{1}{\sigma}\xi_{h,1}5^{1/2}m^{1/2}\left(\log\frac{1}{\zeta}\right)^{1/2}+\sigma\xi_{h,2}5^{3/2}m^{3/2}\left(\log\frac{1}{\zeta}\right)^{3/2}$
$\leq\left(\frac{1}{\sigma}\xi_{h,1}5^{1/2}m^{1/2}+\sigma\xi_{h,2}5^{3/2}m^{3/2}\right)\left(\log\frac{1}{\zeta}\right)^{3/2}$
$=\xi_{h,3}\left(\log\frac{1}{\zeta}\right)^{3/2},$
proving the first claim. As for the second claim, note that using (60), we have
(62) ^$\|\widehat{\nabla}\widetilde{J}_{h}(K)\|^{2}_{F}\leq\frac{1}{\sigma^{2}}\xi_{h,1}^{2}\|\eta_{h}\|^{2}+2\xi_{h,1}\xi_{h,2}\|\eta_{h}\|^{4}+\sigma^{2}\xi_{h,2}^{2}\|\eta_{h}\|^{6}.$
Now since $\|\eta_{h}\|\sim\chi(m)$ whose moments are known, taking an expectation on (62) results in
${\mathbb{E}}\left[\|\widehat{\nabla}\widetilde{J}_{h}(K)\|^{2}_{F}\right]$ $\leq\frac{1}{\sigma^{2}}\xi_{h,1}^{2}{\mathbb{E}}[\|\eta_{\hat{t}}\|^{2}]+2\xi_{h,1}\xi_{h,2}{\mathbb{E}}[\|\eta_{\hat{t}}\|^{4}]+\sigma^{2}\xi_{h,2}^{2}{\mathbb{E}}[\|\eta_{\hat{t}}\|^{6}]$
@@
We next provide some useful properties of the cost function $\widetilde{J}_{h}(K)$ in the following lemma.
@@
where $\widetilde{K}^{*}_{h}$ is the global minimizer of $\widetilde{J}_{h}$. Moreover, assuming that $\|\widetilde{P}_{h+1}-P^{*}_{h+1}\|\leq a$, we have that for all $K_{1},K_{2}\in{\mathbb{R}}^{m\times n}$,
@@
Moreover, let ${\mathcal{F}}_{t}$ denote the $\sigma$-algebra containing the randomness up to iteration $t$ of the inner loop of the algorithm for each $h\in\{0,1,\dotsc,N-1\}$ (including $K_{h,t}$ but not ^$\widehat{\nabla}\widetilde{J}_{h}(K_{h,t})$). We then define
@@
(68) ${\mathbb{E}}[\Delta_{t+1}|{\mathcal{F}}_{t}]1_{\tau>t}\leq\left(\left(1-\mu\alpha_{h,t}\right)\Delta_{t}+\frac{L\alpha_{h,t}^{2}}{2}\xi_{h,4}\right)1_{\tau>t},$
@@
(70) =$\overset{\mathrm{(i)}}{=}-\alpha_{h,t}\|\nabla\widetilde{J}_{h}(K_{h,t})\|_{F}^{2}1_{\tau>t}+\frac{L\alpha_{h,t}^{2}}{2}\xi_{h,4}1_{\tau>t}$
(71) $\overset{\mathrm{(ii)}}{\leq}-\alpha_{h,t}\mu\Delta_{t}1_{\tau>t}+\frac{L\alpha_{h,t}^{2}}{2}\xi_{h,4}1_{\tau>t},$
@@
(72) ${\mathbb{E}}[\Delta_{t+1}|{\mathcal{F}}_{t}]1_{\tau>t}\leq\left(\left(1-\mu\alpha_{h,t}\right)\Delta_{t}+\frac{L\alpha_{h,t}^{2}}{2}\xi_{h,4}\right)1_{\tau>t},$
@@
(73) $\alpha_{h,t}=\frac{2}{\mu}\frac{1}{t+\theta_{h}}\quad\text{for}\quad\theta_{h}=\max\{2,\frac{2L\xi_{h,4}}{\mu^{2}\widetilde{J}_{h}(K_{h,0})}\},$
@@
Moreover, the event $\{\tau\geq T_{h}\}$ happens with probability of at least $\frac{3}{10}\zeta$.
@@
(77) =$\overset{\mathrm{(ii)}}{=}{\mathbb{E}}[{\mathbb{E}}[\Delta_{k+1}|{\mathcal{F}}_{k}]1_{\tau>k}],$
@@
(78) ${\mathbb{E}}[\Delta_{k+1}|{\mathcal{F}}_{k}]1_{\tau>k}$ $\leq\left(\left(1-\mu\alpha_{k}\right)\Delta_{k}+\frac{L\alpha_{k}^{2}}{2}\xi_{h,4}\right)1_{\tau>k}$
(79) =$\overset{\mathrm{(i)}}{=}\left(1-\frac{2}{k+\theta_{h}}\right)\Delta_{k}1_{\tau>k}+\frac{2L\xi_{h,4}}{\mu^{2}}\left(\frac{1}{k+\theta_{h}}\right)^{2},$
@@
(80) ${\mathbb{E}}[\Delta_{k+1}1_{\tau>k+1}]$ $\leq\left(1-\frac{2}{k+\theta_{h}}\right){\mathbb{E}}[\Delta_{k}1_{\tau>k}]+\frac{2L\xi_{h,4}}{\mu^{2}}\left(\frac{1}{k+\theta_{h}}\right)^{2}$
(81) $\overset{\mathrm{(i)}}{\leq}\left(1-\frac{2}{k+\theta_{h}}\right)\frac{\theta_{h}\widetilde{J}_{h}(K_{h,0})}{k+\theta_{h}}+\frac{2L\xi_{h,4}}{\mu^{2}}\left(\frac{1}{k+\theta_{h}}\right)^{2}$
(82) $=\left(1-\frac{1}{k+\theta_{h}}\right)\frac{\theta_{h}\widetilde{J}_{h}(K_{h,0})}{k+\theta_{h}}-\frac{1}{(k+\theta_{h})^{2}}\left(\theta_{h}\widetilde{J}_{h}(K_{h,0})-\frac{2L\xi_{h,4}}{\mu^{2}}\right)$
@@
$\theta_{h}\widetilde{J}_{h}(K_{h,0})-\frac{2L\xi_{h,4}}{\mu^{2}}\geq 0,$
@@
(85) $Y_{t}:=\Delta_{t\wedge\tau}+\frac{4L\xi_{h,4}}{\mu^{2}}\frac{1}{t+\theta_{h}}.$
@@
(86) ${\mathbb{E}}[Y_{t+1}|{\mathcal{F}}_{t}]$ $={\mathbb{E}}[\Delta_{t+1\wedge\tau}|{\mathcal{F}}_{t}]+\frac{4L\xi_{h,4}}{\mu^{2}}\frac{1}{t+\theta_{h}+1}$
(87) $={\mathbb{E}}[\Delta_{t+1\wedge\tau}(1_{\tau\leq t}+1_{\tau>t})|{\mathcal{F}}_{t}]+\frac{4L\xi_{h,4}}{\mu^{2}}\frac{1}{t+\theta_{h}+1}$
(88) $={\mathbb{E}}[\Delta_{t+1\wedge\tau}1_{\tau\leq t}|{\mathcal{F}}_{t}]+{\mathbb{E}}[\Delta_{t+1\wedge\tau}|{\mathcal{F}}_{t}]1_{\tau>t}+\frac{4L\xi_{h,4}}{\mu^{2}}\frac{1}{t+\theta_{h}+1}.$
@@
(90) ${\mathbb{E}}[\Delta_{t+1\wedge\tau}|{\mathcal{F}}_{t}]1_{\tau>t}$ $\overset{\mathrm{(i)}}{\leq}\left(1-\frac{2}{t+\theta_{h}}\right)\Delta_{t}1_{\tau>t}+\frac{2L\xi_{h,4}}{\mu^{2}}\left(\frac{1}{t+\theta_{h}}\right)^{2}1_{\tau>t}$
(91) $\leq\Delta_{t}1_{\tau>t}+\frac{2L\xi_{h,4}}{\mu^{2}}\left(\frac{1}{t+\theta_{h}}\right)^{2},$
@@
(92) ${\mathbb{E}}[Y_{t+1}|{\mathcal{F}}_{t}]$ $\leq\Delta_{t\wedge\tau}1_{\tau\leq t}+\Delta_{t}1_{\tau>t}+\frac{2L\xi_{h,4}}{\mu^{2}}\left(\frac{1}{t+\theta_{h}}\right)^{2}+\frac{4L\xi_{h,4}}{\mu^{2}}\frac{1}{t+\theta_{h}+1}$
(93) $=\Delta_{t\wedge\tau}+\frac{2L\xi_{h,4}}{\mu^{2}}\left(\frac{1}{(t+\theta_{h})^{2}}+\frac{2}{t+\theta_{h}+1}\right)$
(94) $\overset{\mathrm{(i)}}{\leq}\Delta_{t\wedge\tau}+\frac{2L\xi_{h,4}}{\mu^{2}}\left(\frac{2}{t+\theta_{h}}\right)$
@@
${\mathbb{P}}\{\tau\leq T_{h}\}$ =$={\mathbb{P}}\left\{\max_{t\in[T_{h}]}\Delta_{t}>10\zeta^{-1}\widetilde{J}_{h}(K_{h,0})\right\}$
@@
$\overset{\mathrm{(i)}}{\leq}{\mathbb{P}}\left\{\max_{t\in[T_{h}]}Y_{t}\geq 10\zeta^{-1}\widetilde{J}_{h}(K_{h,0})\right\},$
@@
${\mathbb{P}}\{\tau\leq T_{h}\}$ ${\leq}\frac{\zeta{\mathbb{E}}[Y_{0}]}{10\widetilde{J}_{h}(K_{h,0})}=\frac{\zeta\left(\Delta_{0}+\frac{4L\xi_{h,4}}{\mu^{2}}\frac{1}{\theta_{h}}\right)}{10\widetilde{J}_{h}(K_{h,0})}.$

### v3

@@ 4. The Inner Loop and Policy Gradient
$Q_{h}(x_{0},-Kx_{0}+\sigma\eta)$
$=$ $x_{0}^{\top}Qx_{0}+(-Kx_{0}+\sigma\eta)^{\top}R(-Kx_{0}+\sigma\eta)+(Ax_{0}+B(-Kx_{0}+\sigma\eta))^{\top}\widetilde{P}_{h+1}(Ax_{0}+B(-Kx_{0}+\sigma\eta))$
$\leq$ $\|Q\|C_{m}+\|R\|\|-Kx_{0}+\sigma\eta\|^{2}+\|\widetilde{P}_{h+1}\|\|Ax_{0}+B(-Kx_{0}+\sigma\eta)\|^{2}.$
@@
(52) $Q_{h}(x_{0},-Kx_{0}+\sigma\eta)$
(53) $\leq$ $\|Q\|C_{m}+2\|R\|(\widetilde{C}_{h}C_{m}+\sigma^{2}\|\eta\|^{2})+2(\|Q_{N}\|+a)\|A\|^{2}C_{m}+4(\|Q_{N}\|+a)\|B\|^{2}(\widetilde{C}_{h}C_{m}+\sigma^{2}\|\eta\|^{2}))$
(54) $=$ $C_{m}(\|Q\|+2\|R\|\widetilde{C}_{h})+2C_{m}(\|Q_{N}\|+a)(\|A\|^{2}+2\|B\|^{2}\widetilde{C}_{h})+2\left(\|R\|+2(\|Q_{N}\|+a)\|B\|^{2}\right)\sigma^{2}\|\eta\|^{2},$
@@
(55) $\|\widehat{\nabla}\widetilde{J}_{h}(K)\|_{F}$ $\leq\frac{1}{\sigma}\Bigg(\|Q\|+2\|R\|\widetilde{C}_{h}+2(\|Q_{N}\|+a)(\|A\|^{2}+2\|B\|^{2}\widetilde{C}_{h})\Bigg)C_{m}^{3/2}\|\eta\|$
(56) $\quad+2\left(\|R\|+2(\|Q_{N}\|+a)\|B\|^{2}\right)\sigma C_{m}^{1/2}\|\eta\|^{3}$
(57) $=\frac{1}{\sigma}\xi_{h,1}\|\eta\|+\sigma\xi_{h,2}\|\eta\|^{3}.$
Now note that using (57), we have
(58) $\|\widehat{\nabla}\widetilde{J}_{h}(K)\|^{2}_{F}\leq\frac{1}{\sigma^{2}}\xi_{h,1}^{2}\|\eta\|^{2}+2\xi_{h,1}\xi_{h,2}\|\eta\|^{4}+\sigma^{2}\xi_{h,2}^{2}\|\eta\|^{6}.$
Now since $\|\eta\|\sim\chi(m)$ whose moments are known, taking an expectation on (58) results in
${\mathbb{E}}\left[\|\widehat{\nabla}\widetilde{J}_{h}(K)\|^{2}_{F}\right]$ $\leq\frac{1}{\sigma^{2}}\xi_{h,1}^{2}{\mathbb{E}}[\|\eta\|^{2}]+2\xi_{h,1}\xi_{h,2}{\mathbb{E}}[\|\eta\|^{4}]+\sigma^{2}\xi_{h,2}^{2}{\mathbb{E}}[\|\eta\|^{6}]$
@@
We next establish some useful properties of the cost function $\widetilde{J}_{h}(K)$ in the following lemma.
@@
where $\widetilde{K}^{*}_{h}$ is the global minimizer of $\widetilde{J}_{h}$. Moreover, assuming that $\|\widetilde{P}_{h+1}-P^{*}_{h+1}\|\leq a$, we have that for all $K_{1},K_{2}\in^{m\times n}$,
@@
Moreover, let ${\mathcal{F}}_{t}$ denote the $\sigma$-algebra containing the randomness up to iteration $t$ of the inner loop of the algorithm for each $h\in\{0,1,\dotsc,N-1\}$ (including $K_{h,t}$ but not $\widehat{\nabla}\widetilde{J}_{h}(K_{h,t})$). We then define
@@
Combining the unbiasedness of the estimate (Proposition 4.1), the second-moment bound on its size (Lemma 4.1), and the regularity properties of $\widetilde{J}_{h}$ (Lemma 4.2), we now show the following descent recursion for the optimality gap:
@@
(64) ${\mathbb{E}}[\Delta_{t+1}|{\mathcal{F}}_{t}]1_{\tau>t}\leq\left(\left(1-\mu\alpha_{h,t}\right)\Delta_{t}+\frac{L\alpha_{h,t}^{2}}{2}\xi_{h,3}\right)1_{\tau>t},$
@@
(66) $\overset{\mathrm{(i)}}{\leq}-\alpha_{h,t}\|\nabla\widetilde{J}_{h}(K_{h,t})\|_{F}^{2}1_{\tau>t}+\frac{L\alpha_{h,t}^{2}}{2}\xi_{h,3}1_{\tau>t}$
(67) $\overset{\mathrm{(ii)}}{\leq}-\alpha_{h,t}\mu\Delta_{t}1_{\tau>t}+\frac{L\alpha_{h,t}^{2}}{2}\xi_{h,3}1_{\tau>t},$
@@
(68) ${\mathbb{E}}[\Delta_{t+1}|{\mathcal{F}}_{t}]1_{\tau>t}\leq\left(\left(1-\mu\alpha_{h,t}\right)\Delta_{t}+\frac{L\alpha_{h,t}^{2}}{2}\xi_{h,3}\right)1_{\tau>t},$
@@
(69) $\alpha_{h,t}=\frac{2}{\mu}\frac{1}{t+\theta_{h}}\quad\text{for}\quad\theta_{h}=\max\{2,\frac{2L\xi_{h,3}}{\mu^{2}\widetilde{J}_{h}(K_{h,0})}\},$
@@
Moreover, the event $\{\tau\leq T_{h}\}$ happens with probability of at most $\frac{3}{10}\zeta$.
@@
(73) $\overset{\mathrm{(ii)}}{=}{\mathbb{E}}[{\mathbb{E}}[\Delta_{k+1}|{\mathcal{F}}_{k}]1_{\tau>k}],$
@@
(74) ${\mathbb{E}}[\Delta_{k+1}|{\mathcal{F}}_{k}]1_{\tau>k}$ $\leq\left(\left(1-\mu\alpha_{k}\right)\Delta_{k}+\frac{L\alpha_{k}^{2}}{2}\xi_{h,3}\right)1_{\tau>k}$
(75) $\overset{\mathrm{(i)}}{=}\left(1-\frac{2}{k+\theta_{h}}\right)\Delta_{k}1_{\tau>k}+\frac{2L\xi_{h,3}}{\mu^{2}}\left(\frac{1}{k+\theta_{h}}\right)^{2},$
@@
(76) ${\mathbb{E}}[\Delta_{k+1}1_{\tau>k+1}]$ $\leq\left(1-\frac{2}{k+\theta_{h}}\right){\mathbb{E}}[\Delta_{k}1_{\tau>k}]+\frac{2L\xi_{h,3}}{\mu^{2}}\left(\frac{1}{k+\theta_{h}}\right)^{2}$
(77) $\overset{\mathrm{(i)}}{\leq}\left(1-\frac{2}{k+\theta_{h}}\right)\frac{\theta_{h}\widetilde{J}_{h}(K_{h,0})}{k+\theta_{h}}+\frac{2L\xi_{h,3}}{\mu^{2}}\left(\frac{1}{k+\theta_{h}}\right)^{2}$
(78) $=\left(1-\frac{1}{k+\theta_{h}}\right)\frac{\theta_{h}\widetilde{J}_{h}(K_{h,0})}{k+\theta_{h}}-\frac{1}{(k+\theta_{h})^{2}}\left(\theta_{h}\widetilde{J}_{h}(K_{h,0})-\frac{2L\xi_{h,3}}{\mu^{2}}\right)$
@@
$\theta_{h}\widetilde{J}_{h}(K_{h,0})-\frac{2L\xi_{h,3}}{\mu^{2}}\geq 0,$
@@
(81) $Y_{t}:=\Delta_{t\wedge\tau}+\frac{4L\xi_{h,3}}{\mu^{2}}\frac{1}{t+\theta_{h}}.$
@@
(82) ${\mathbb{E}}[Y_{t+1}|{\mathcal{F}}_{t}]$ $={\mathbb{E}}[\Delta_{t+1\wedge\tau}|{\mathcal{F}}_{t}]+\frac{4L\xi_{h,3}}{\mu^{2}}\frac{1}{t+\theta_{h}+1}$
(83) $={\mathbb{E}}[\Delta_{t+1\wedge\tau}(1_{\tau\leq t}+1_{\tau>t})|{\mathcal{F}}_{t}]+\frac{4L\xi_{h,3}}{\mu^{2}}\frac{1}{t+\theta_{h}+1}$
(84) $={\mathbb{E}}[\Delta_{t+1\wedge\tau}1_{\tau\leq t}|{\mathcal{F}}_{t}]+{\mathbb{E}}[\Delta_{t+1\wedge\tau}|{\mathcal{F}}_{t}]1_{\tau>t}+\frac{4L\xi_{h,3}}{\mu^{2}}\frac{1}{t+\theta_{h}+1}.$
@@
(86) ${\mathbb{E}}[\Delta_{t+1\wedge\tau}|{\mathcal{F}}_{t}]1_{\tau>t}$ $\overset{\mathrm{(i)}}{\leq}\left(1-\frac{2}{t+\theta_{h}}\right)\Delta_{t}1_{\tau>t}+\frac{2L\xi_{h,3}}{\mu^{2}}\left(\frac{1}{t+\theta_{h}}\right)^{2}1_{\tau>t}$
(87) $\leq\Delta_{t}1_{\tau>t}+\frac{2L\xi_{h,3}}{\mu^{2}}\left(\frac{1}{t+\theta_{h}}\right)^{2},$
@@
(88) ${\mathbb{E}}[Y_{t+1}|{\mathcal{F}}_{t}]$ $\leq\Delta_{t\wedge\tau}1_{\tau\leq t}+\Delta_{t}1_{\tau>t}+\frac{2L\xi_{h,3}}{\mu^{2}}\left(\frac{1}{t+\theta_{h}}\right)^{2}+\frac{4L\xi_{h,3}}{\mu^{2}}\frac{1}{t+\theta_{h}+1}$
(89) $=\Delta_{t\wedge\tau}+\frac{2L\xi_{h,3}}{\mu^{2}}\left(\frac{1}{(t+\theta_{h})^{2}}+\frac{2}{t+\theta_{h}+1}\right)$
(90) $\overset{\mathrm{(i)}}{\leq}\Delta_{t\wedge\tau}+\frac{2L\xi_{h,3}}{\mu^{2}}\left(\frac{2}{t+\theta_{h}}\right)$
@@
${\mathbb{P}}\{\tau\leq T_{h}\}$ $={\mathbb{P}}\left\{\max_{t\in[T_{h}]}\Delta_{t}>10\zeta^{-1}\widetilde{J}_{h}(K_{h,0})\right\}$
@@
$\overset{\mathrm{(i)}}{\leq}{\mathbb{P}}\left\{\max_{t\in[T_{h}]}Y_{t}\geq 10\zeta^{-1}\widetilde{J}_{h}(K_{h,0})\right\}$
@@
${\mathbb{P}}\{\tau\leq T_{h}\}$ ${\leq}\frac{\zeta{\mathbb{E}}[Y_{0}]}{10\widetilde{J}_{h}(K_{h,0})}$
$=\frac{\zeta\left(\Delta_{0}+\frac{4L\xi_{h,3}}{\mu^{2}}\frac{1}{\theta_{h}}\right)}{10\widetilde{J}_{h}(K_{h,0})}.$